OpenStudy (anonymous):

The ages of three brothers are consecutive even integers. Three times the age of the youngest brother exceeds the oldest brother’s age by 48 years. What is the age of the youngest brother?

OpenStudy (anonymous):

x - 2, x, x + 2

OpenStudy (anonymous):

3 (x - 2) = (x + 2) + 48 x = 28 {26, 28, 30}

OpenStudy (anonymous):

Youngest brother's age is 26 years old.